OpenStudy (anonymous):

bz, cx and ay are heights of the equilateral triangle abc drawn above. If from b down to intersecting point of the heights is 4 square root 3, what is the area of the triangle abc?

OpenStudy (anonymous):

side of equiltera triangle = 2/sqrt(3) * height of equilateral triangle area of equilateral triangle =sqrt(3)/4 * (side of equilateral triangle )^2

OpenStudy (anonymous):

here height of equilateral triangle =4*sqrt(3) hence first calculate side and then calculate are using the fromula baove

OpenStudy (anonymous):

If from b down to intersecting point of the heights is 4 square root 3, so here the height is not full, but till the point, so is 4*sqrt(3) half of the height???

OpenStudy (anonymous):

side of equiltera triangle = 8 8^2*sqrt(3)/4= 16sqrt(3)

OpenStudy (anonymous):

oops i didn't read the question completely and assumed the height as 4*sqrt(3) for equilateral triangle the heights intersect at the centroid . centroid here divides the heights in the ratio of 2:1 (in general for any triangles centroid always divides medians in the ratio of 2:1) so here 2/3 * height = 4*sqrt(3) hence height =3/2 * 4*sqrt(3)